# Model Merging in the Era of Large Language Models: Methods, Applications, and Future Directions

### 10 Mar 2026 v1

@@ Abstract
Model merging has emerged as a transformative paradigm for combining the capabilities of multiple neural networks into a single unified model without additional training. With the rapid proliferation of fine-tuned large language models (LLMs), merging techniques offer a computationally efficient alternative to ensembles and full retraining, enabling practitioners to compose specialized capabilities at minimal cost. This survey presents a comprehensive and structured examination of model merging in the LLM era through the FUSE taxonomy, a four-dimensional framework organized along Foundations, Unification Strategies, Scenarios, and Ecosystem. We first establish the theoretical underpinnings of merging, including loss landscape geometry, mode connectivity, and the linear mode connectivity hypothesis. We then systematically review the algorithmic landscape, spanning weight averaging, task vector arithmetic, sparsification-enhanced methods, mixture-of-experts architectures, and evolutionary optimization approaches. For each method family, we analyze the core formulation, highlight representative works, and discuss practical trade-offs. We further examine downstream applications across multi-task learning, safety alignment, domain specialization, multilingual transfer, and federated learning. Finally, we survey the supporting ecosystem of open-source tools, community platforms, and evaluation benchmarks, and identify key open challenges including theoretical gaps, scalability barriers, and standardization needs. This survey aims to equip researchers and practitioners with a structured foundation for advancing model merging.
@@ 1 Introduction
Model merging combines the parameters of multiple trained neural networks into a single unified model, inheriting their capabilities without inference-time ensemble overhead (Li et al., 2023a). Unlike ensembles that aggregate predictions from separate models at runtime, merging operates directly in weight space, yielding a single model whose cost matches any individual source. Early work showed that averaging parameters of models trained with different random seeds improves generalization performance (Izmailov et al., 2018). In the early days, these techniques were primarily applied to homogeneous models trained on identical datasets, with Stochastic Weight Averaging emerging as a principled approach for combining checkpoints along a single optimization trajectory (Garipov et al., 2018). With the advent of large-scale pretrained language model (Devlin et al., 2019)s, the scope of model merging expanded dramatically, enabling the combination of models fine-tuned for entirely different tasks into unified systems exhibiting multi-task capabilities (Ilharco et al., 2023). This evolution from checkpoint averaging to cross-task composition marks a shift in how the community conceptualizes neural network knowledge reuse.
@@
The emergence of pretrained foundation models such as LLaMA (Touvron et al., 2023a), Mistral (Jiang et al., 2023), Qwen (Hu et al., 2023), DeepSeek (Bi et al., 2024), Gemma (Team, 2024a), and CLIP (Radford et al., 2021) has created conditions that make model merging both practically viable and scientifically compelling. The key insight is that models fine-tuned from a shared pretrained initialization tend to reside within the same loss basin, exhibiting linear mode connectivity (Singh and Jaggi, 2020) that permits direct weight interpolation without traversing high-loss barriers (Garipov et al., 2018). This theoretical grounding, combined with the proliferation of openly available fine-tuned large language models, has sparked substantial interest in merging methodologies. The past two years have witnessed an explosion of research activity, with novel techniques such as task vector arithmetic, sparsification-enhanced merging, and activation-informed (Hammoud et al., 2024) approaches achieving notable success in combining specialized models (Yadav et al., 2023; Yu et al., 2023; Huang et al., 2024). Practically, merged models have achieved top rankings on competitive leaderboards including the Open LLM Leaderboard (Li et al., 2023b), demonstrating that strategic model combination can yield capabilities exceeding those of individual fine-tuned variants (Wortsman et al., 2022a; Jang et al., 2024). The establishment of dedicated toolkits such as mergekit (Goddard et al., 2024) has further democratized access to these techniques, enabling experiment with sophisticated merging strategies without extensive technical expertise. This confluence of theoretical understanding, empirical success, and practical tooling (Goddard et al., 2024) has positioned model merging as a central technique in modern LLM development.
@@
Several recent surveys have begun to address aspects of this rapidly evolving landscape. Yang et al. (2024b) and Li et al. (2023a) provide a broad treatment of model fusion techniques across deep learning, while Yadav et al. (2024a) focus specifically on mixture-of-experts style composition approaches. Yang et al. (2024b) examine collaborative strategies more broadly, encompassing both merging and ensemble methods. However, existing surveys lack a unified treatment that systematically connects theoretical foundations explaining why merging succeeds, algorithmic strategies for performing merging effectively, downstream application scenarios where merging provides value, and the practical ecosystem supporting deployment. No prior work has comprehensively addressed the full spectrum from loss landscape geometry and mode connectivity theory through to evaluation methodologies and community tooling, nor proposed a unified taxonomy for navigating this space.
This survey provides a systematic treatment of model merging in the LLM era, bridging theoretical understanding with practical application. Table 1 summarizes the major open-weight foundation models (Brown et al., 2020; Team, 2024b) commonly used in merging research. As shown in Figure 2, our survey is organized around the FUSE taxonomy, a four-dimensional framework encompassing Foundations, Unification Strategies, Scenarios, and Ecosystem, that structures the field according to why merging works, how merging is performed, where merging provides value, and what resources support merging in practice. Figure 1 provides a schematic overview of the model merging pipeline. The main contributions of this survey are as follows: (1) We propose the FUSE taxonomy as a principled organizational framework that comprehensively categorizes model merging research across theoretical foundations including loss landscape geometry, mode connectivity, and weight space symmetries; algorithmic unification strategies spanning weight-space averaging, task vector arithmetic, geometric interpolation, and architecture-level composition; application scenarios including capability augmentation, alignment and safety, and efficiency-driven deployment; and ecosystem infrastructure encompassing toolkits, benchmark s, and community resources. (2) We provide an in-depth technical analysis of merging methodologies, elucidating the mathematical principles underlying each approach and offering comparative insights into their respective strengths, limitations, and appropriate use cases. (3) We conduct a systematic examination of application domains, synthesizing empirical evidence regarding when and how model merging delivers practical benefits across multi-task generalization, multilingual transfer, preference alignment, and federated learning scenarios. (4) We identify open challenges and articulate promising future research (Li et al., 2025) directions, including theoretical gaps in understanding mergeability, scalability limitations for increasingly large models, and opportunities for automated and dynamic merging systems. The remainder of this survey proceeds as follows: Section 2 establishes theoretical foundations; Sections 3–5 present unification strategies with increasing sophistication; Section 6 examines application scenarios; and Section 7 discusses the ecosystem, open challenges, and future directions.
@@
Table 1: Overview of major open-weight foundation models (Brown et al., 2020; Team, 2024b) commonly used in model merging research. These models serve as the base for fine-tuning and subsequent merging experiments.
@@
Gemma (Team, 2024a) Google 2B–7B 2024 Decoder-only RoPE, Multi-Query Attn
@@
Figure 1: Overview of the model merging pipeline. Task-specific models $\{\theta_{1},\dots,\theta_{K}\}$, obtained by fine-tuning a shared pretrained checkpoint $\theta_{0}$, are combined via one of several strategies (Sections 3–5) into a single model $\theta_{\mathrm{merged}}$, which is then evaluated on all source tasks. Gray vertical bars represent the many-to-many mapping between model inputs and merging strategies.
@@ 2 Theoretical Foundations of Model Merging
Why does model merging work? We examine three interconnected pillars: the geometric structure of loss landscapes that permits effective weight interpolation (Ren et al., 2025), the phenomenon of mode connectivity enabling low-loss paths between independently trained solutions (Draxler et al., 2018; Lubana et al., 2023), and weight-space symmetries, particularly permutation invariance (Jordan et al., 2022; Entezari et al., 2022), that must be resolved for effective merging (Ainsworth et al., 2023).
@@ 2.1 Loss Landscape Geometry and Convexity Properties in Neural Networks
The success of model merging fundamentally depends on the geometric properties of the loss landscape in which neural network parameters reside. Understanding this landscape, particularly how loss values vary as one traverses the high-dimensional parameter space, provides essential insight into why weight interpolation between independently trained models can yield functional merged networks rather than catastrophic performance degradation.
@@
Neural network loss landscapes exhibit remarkably complex topologies characterized by numerous local minima, saddle points, and regions of varying curvature. However, a crucial observation that underpins model merging is that modern deep networks, particularly when sufficiently overparameterized, display favorable geometric properties that deviate significantly from the pathological non-convex landscapes one might naively expect. Singh et al. (2024) provide a systematization of these landscape properties, demonstrating that overparameterization fundamentally alters the geometry of the optimization surface. Specifically, when parameters substantially outnumber the minimum required for the task, the loss landscape exhibits large connected regions of near-optimal solutions rather than isolated minima.
@@
The concept of basin characteristics, encompassing both width and flatness, plays a central role in determining whether weight interpolation between two solutions will preserve model functionality. A loss basin can be formally characterized by considering the set of parameters $\mathcal{B}_{\epsilon}(\theta^{*})=\{\theta:\mathcal{L}(\theta)<\mathcal{L}(\theta^{*})+\epsilon\}$ surrounding a local minimum $\theta^{*}$. The width of this basin, often measured through the eigenspectrum of the Hessian matrix $\nabla^{2}\mathcal{L}(\theta^{*})$, determines how much perturbations from the optimum increase loss. Wide, flat basins permit substantial movement in parameter space without significant performance degradation, directly enabling weight averaging.
Garipov et al. (2018) investigated the empirical phenomenon of monotonic linear interpolation, wherein the loss along the linear path $\theta_{\alpha}=(1-\alpha)\theta_{1}+\alpha\theta_{2}$ between two trained networks exhibits monotonically decreasing behavior toward the interpolated midpoint under certain conditions. This monotonicity property suggests that linear interpolation not only avoids catastrophic loss barriers but can actually discover solutions with lower loss than either endpoint, a finding with important implications for model merging strategies that rely on simple averaging operations.
@@
The relationship between loss landscape geometry and generalization performance introduces additional considerations for model merging. Flat minima, characterized by low curvature in the Hessian eigenspectrum, have been associated with improved generalization properties  (Izmailov et al., 2018). When merging models that reside in flat regions of the loss landscape, the resulting averaged parameters are more likely to remain within a low-loss region, thereby preserving generalization capabilities. Conversely, sharp minima (Keskar et al., 2017) with high curvature present challenges for merging, as even small perturbations induced by averaging can push parameters into high-loss regions.
Lubana et al. (2023) further examined the dynamics of weight interpolation, revealing that interpolated models can exhibit transient performance characteristics that differ from simple averaging predictions. Their analysis demonstrates that the geometry of the loss landscape along interpolation paths depends critically on training dynamics and the specific optimization trajectory taken to reach each solution.
@@
Mathematical analysis of overparameterized networks reveals that the loss surface increasingly resembles a high-dimensional valley structure as model capacity grows. Let $\mathcal{M}$ denote the manifold of global minima; in the overparameterized regime, $\mathcal{M}$ becomes a connected, often convex subspace (Tang et al., 2024d) within which arbitrary linear combinations of solutions remain optimal. This geometric insight, explored in (Ferbach et al., 2024), provides theoretical justification for why weight averaging succeeds when models share sufficient structural commonality, their solutions lie within the same connected low-loss region, rendering interpolation a geometrically valid operation (Neyshabur et al., 2020).
@@ 2.2 Linear Mode Connectivity Theory and Basin Characteristics
The geometric insight that successful merging requires solutions to occupy connected low-loss regions naturally leads to the formal framework of mode connectivity, which provides a fundamental theoretical lens through which to understand when and why model merging succeeds. Mode connectivity refers to the existence of paths in weight space connecting different trained solutions along which the loss remains consistently low. This concept has emerged as central to explaining the empirical success of weight interpolation methods, as thoroughly surveyed in (Yang et al., 2024b).
@@
Formally, two neural network solutions $\theta_{1}$ and $\theta_{2}$ are said to exhibit linear mode connectivity (Singh and Jaggi, 2020) if the loss along the linear interpolation path $\theta(\alpha)=(1-\alpha)\theta_{1}+\alpha\theta_{2}$ for $\alpha\in[0,1]$ satisfies $\mathcal{L}(\theta(\alpha))\leq\max(\mathcal{L}(\theta_{1}),\mathcal{L}(\theta_{2}))+\epsilon$ for some small tolerance $\epsilon\geq 0$. When $\epsilon=0$, the models reside in the same loss basin with no interpolation barrier. More generally, nonlinear mode connectivity (Singh and Jaggi, 2020) relaxes the path constraint to allow curved trajectories, typically parameterized as Bézier curves or piecewise linear paths. This distinction has practical implications: linear mode connectivity (Singh and Jaggi, 2020) enables direct weight averaging, while nonlinear connectivity may require more sophisticated merging procedures.
@@
The analysis of loss barriers, the maximum loss increase encountered along interpolation paths, provides quantitative measures for characterizing mode connectivity strength. Singh et al. (2024) present a systematization of methods for computing and interpreting these barriers, noting that barrier height correlates inversely with merging success. Specifically, given solutions $\theta_{1}$ and $\theta_{2}$, the loss barrier is defined as:
$\Delta\mathcal{L}=\max_{\alpha\in[0,1]}\mathcal{L}(\theta(\alpha))-\frac{1}{2}(\mathcal{L}(\theta_{1})+\mathcal{L}(\theta_{2}))$ (1)
@@
When $\Delta\mathcal{L}\leq\epsilon$ for some small threshold $\epsilon$, we say the models exhibit $\epsilon$-linear mode connectivity (Singh and Jaggi, 2020). The relationship between loss barrier magnitude and merging success can be formally characterized through the following theorem:
Proposition 1 (Merging Error Bound (Singh et al., 2024)).
@@
Let $\theta_{1},\theta_{2}$ be two models with shared pretrained initialization $\theta_{0}$, and let $\theta_{\text{avg}}=\frac{1}{2}(\theta_{1}+\theta_{2})$. Under $L$-smoothness of the loss function $\mathcal{L}$, the performance degradation of the averaged model is bounded by:
$\mathcal{L}(\theta_{\text{avg}})-\frac{1}{2}(\mathcal{L}(\theta_{1})+\mathcal{L}(\theta_{2}))\leq\frac{L}{8}\|\theta_{1}-\theta_{2}\|^{2}$ (2)
@@
This bound reveals why shared initialization is critical: models fine-tuned from the same pretrained checkpoint exhibit smaller $\|\theta_{1}-\theta_{2}\|$, yielding tighter merging guarantees.
@@
A critical insight enabling practical model merging in the large language model era concerns the role of shared pretrained initialization. When multiple models are fine-tuned from a common pretrained checkpoint, they exhibit substantially stronger linear mode connectivity (Singh and Jaggi, 2020) compared to independently trained models (Frankle et al., 2020; Garipov et al., 2018). This phenomenon arises because fine-tuning typically induces relatively small parameter perturbations that remain within the same loss basin as the pretrained model, whereas training from random initialization explores fundamentally different regions of the loss landscape.
@@
Empirical investigations have employed various path parameterizations to probe connectivity structure. Quadratic Bézier curves, parameterized as $\theta(\alpha)=(1-\alpha)^{2}\theta_{1}+2\alpha(1-\alpha)\theta_{m}+\alpha^{2}\theta_{2}$ with learnable midpoint $\theta_{m}$, can discover low-loss nonlinear paths even when linear barriers exist. Ren et al. (2025) demonstrated that trained networks frequently exhibit monotonic interpolation properties, where loss decreases steadily from one endpoint toward the other, suggesting asymmetric basin structures.
@@
The work of Nobari et al. (2025) revealed that apparent loss barriers often arise from feature distribution mismatches rather than genuine geometric separation, proposing activation renormalization to restore connectivity. This finding suggests that many observed failures of linear mode connectivity (Singh and Jaggi, 2020) stem from correctable statistical artifacts rather than fundamental incompatibility between solutions.
@@
Understanding basin characteristics extends beyond barrier analysis to encompass basin width and shape. Lubana et al. (2023) demonstrated that interpolated models can exhibit performance dynamics reflecting the local basin geometry, with wider basins supporting more robust interpolation. These theoretical insights directly inform practical merging strategies by establishing that shared pretraining serves as a powerful mechanism for ensuring models occupy connected regions amenable to linear combination. However, even when models theoretically reside in connected basins, the success of direct weight averaging depends critically on how the models are represented, a consideration that leads naturally to examining the fundamental symmetries inherent in neural network weight spaces.

### 30 Mar 2026 v2

# Model Merging in the Era of Large Language Models Methods, Applications, and Future Directions

@@ Abstract
Model merging combines the parameters of multiple neural networks into a single model without additional training. As fine-tuned large language models (LLMs) proliferate, merging offers a computationally efficient alternative to ensembles and full retraining, enabling practitioners to compose specialized capabilities at minimal cost. This survey examines model merging in the LLM era through the FUSE taxonomy, organized along Foundations, Unification Strategies, Scenarios, and Ecosystem. We first establish the theoretical underpinnings of merging, including loss landscape geometry and mode connectivity, then systematically review the algorithmic space spanning weight averaging, task vector arithmetic, sparsification-enhanced methods, mixture-of-experts architectures, and evolutionary optimization. We further examine downstream applications across multi-task learning, safety alignment, domain specialization, and federated learning, and survey the supporting ecosystem of tools and evaluation benchmarks. Finally, we identify key open challenges and future directions, aiming to equip researchers and practitioners with a structured foundation for advancing model merging.
@@ 1 Introduction
Model merging combines the parameters of multiple trained neural networks into a single unified model that inherits their respective capabilities without the inference-time overhead of ensembles. Whereas ensembles aggregate predictions from separate models at runtime, merging operates directly in weight space, producing a single model whose computational cost matches that of any individual constituent. The idea traces back to the discovery that independently trained networks are connected by low-loss curves in weight space (Garipov et al., 2018) and to Stochastic Weight Averaging (Izmailov et al., 2018), which showed that averaging checkpoints along an optimization trajectory yields flatter, better-generalizing solutions. Large-scale pretrained language models (Devlin et al., 2019) broadened the scope dramatically; models fine-tuned for entirely different tasks could now be merged into unified systems with multi-task capabilities (Ilharco et al., 2023), turning what had been a checkpoint-averaging heuristic into a principled approach to neural network knowledge composition.
@@
Pretrained foundation models such as LLaMA (Touvron et al., 2023a), Mistral (Jiang et al., 2023), Qwen (Hu et al., 2023; Team, 2025), DeepSeek (Bi et al., 2024), Gemma (Team, 2024a), and CLIP (Radford et al., 2021) have created conditions that make model merging both practically viable and scientifically compelling. The key insight is that models fine-tuned from a shared pretrained initialization tend to reside within the same loss basin, exhibiting linear mode connectivity (Frankle et al., 2020) that permits direct weight interpolation without traversing high-loss barriers (Garipov et al., 2018). This theoretical grounding, combined with the proliferation of openly available fine-tuned large language models, has sparked broad interest in merging methodologies.
Rapid methodological progress over the past two years has transformed model merging from a niche technique into a mainstream component of LLM development. Task vector arithmetic (Ilharco et al., 2023), sparsification-enhanced methods such as TIES-Merging (Yadav et al., 2023) and DARE (Yu et al., 2024), and activation-informed approaches (Nobari et al., 2025) now enable the principled combination of specialized models (Huang et al., 2024a). Merged models have reached top rankings on competitive leaderboards including the Open LLM Leaderboard, with strategic model combination yielding capabilities exceeding those of individual fine-tuned variants (Wortsman et al., 2022a; Jang et al., 2024b). Dedicated toolkits such as MergeKit (Goddard et al., 2024) have further democratized access, putting advanced merging strategies within reach of practitioners without deep algorithmic expertise.
@@
Several recent surveys have begun to address aspects of this rapidly evolving field. Yang et al. (2024b) provide a broad treatment of model merging techniques, while Yadav et al. (2025a) focus specifically on mixture-of-experts style composition. Concurrent works include Ruan et al. (2025), who review model merging from task-specific models to unified systems, and Lotfi et al. (2026), who systematically analyze trade-offs among ensembling, merging, and routing for parameter-efficient experts. However, no existing survey provides a unified treatment that simultaneously connects theoretical foundations explaining why merging succeeds, algorithmic strategies for how to merge effectively, downstream scenarios illustrating where merging provides value, and the practical ecosystem supporting deployment. In particular, the full spectrum from loss surface geometry and mode connectivity theory through evaluation methodologies and community tooling has not been covered within a single coherent taxonomic framework.
To fill this gap, this survey provides a structured treatment of model merging in the LLM era, bridging theoretical understanding with practical application. Table 1 summarizes the major open-weight foundation models commonly used in merging research. As shown in Figure 2, we organize the field through the FUSE taxonomy, a four-dimensional scheme encompassing Foundations (why merging works), Unification Strategies (how merging is performed), Scenarios (where merging provides value), and Ecosystem (what resources support merging). Figure 1 provides a schematic overview of the model merging pipeline.
The main contributions are as follows.
  1. 1.
    We propose the FUSE taxonomy as a coherent organizational structure that categorizes model merging research across theoretical foundations, algorithmic strategies, application scenarios, and ecosystem infrastructure.
  2. 2.
    We provide an in-depth technical analysis of merging methodologies, explaining the mathematical principles underlying each approach and offering comparative insights into their strengths and limitations.
  3. 3.
    We conduct a systematic examination of application domains, synthesizing empirical evidence on when and how merging delivers practical benefits across multi-task generalization, multilingual transfer, preference alignment, and federated learning.
  4. 4.
    We identify open challenges and promising future research directions (Li et al., 2025d), including theoretical gaps in understanding mergeability, scalability limitations, and opportunities for automated merging systems.
The remainder of this survey proceeds as follows. Section 2 establishes theoretical foundations. Sections 35 present unification strategies with increasing sophistication. Section 6 examines application scenarios. Section 7 discusses the ecosystem, systematic failure modes (Section 7.4), open challenges, and future directions.
@@
Table 1: Overview of major open-weight foundation models commonly used in model merging research. These models serve as the base for fine-tuning and subsequent merging experiments.
@@
Qwen3 (Team, 2025) Alibaba 0.6B–235B 2025 Dense + MoE Thinking mode, 32K–128K context
@@
Gemma (Team, 2024a) Google 2B–7B 2024 Decoder-only RoPE, GeGLU, MQA (2B)
@@
Figure 1: Overview of the model merging pipeline. Task-specific models $\{\theta_{1},\dots,\theta_{K}\}$, obtained by fine-tuning a shared pretrained checkpoint $\theta_{0}$, are combined via one of several strategies (Sections 3–5) into a single model $\theta_{\text{merged}}$, which is then evaluated on all source tasks. Gray vertical bars represent the many-to-many mapping between model inputs and merging strategies.
@@ 2 Theoretical Foundations of Model Merging
Why does model merging work? We examine three interconnected pillars. First, the geometric structure of loss landscapes that permits effective weight interpolation (Ren et al., 2025), the phenomenon of mode connectivity that permits low-loss paths between independently trained solutions (Draxler et al., 2018; Lubana et al., 2023), and weight-space symmetries, particularly permutation invariance (Jordan et al., 2023; Entezari et al., 2022), that must be resolved for effective merging (Ainsworth et al., 2023).
@@ 2.1 Loss Landscape Geometry and Convexity Properties in Neural Networks
Successful model merging critically depends on the geometric properties of the loss surface. Neural network loss surfaces exhibit complex topologies with numerous local minima, saddle points, and regions of varying curvature. Yet a crucial observation underpinning model merging is that modern overparameterized networks display favorable geometric properties, deviating markedly from the pathological non-convex landscapes one might naively expect. Singh et al. (2024) systematize these properties, showing that when parameters vastly outnumber the minimum required for the task, the loss surface exhibits large connected regions of near-optimal solutions rather than isolated minima.
@@
Basin characteristics, encompassing both width and flatness, play a central role in determining whether weight interpolation between two solutions will preserve model functionality. A loss basin can be formally characterized by considering the set of parameters $\mathcal{B}_{\epsilon}(\theta^{*})=\{\theta:\mathcal{L}(\theta)<\mathcal{L}(\theta^{*})+\epsilon\}$ surrounding a local minimum $\theta^{*}$. The width of this basin, often measured through the eigenspectrum of the Hessian matrix $\nabla^{2}\mathcal{L}(\theta^{*})$, determines how much perturbations from the optimum increase loss. Wide, flat basins permit considerable movement in parameter space without appreciable performance degradation, directly supporting weight averaging.
Garipov et al. (2018) discovered that independently trained neural networks can be connected by low-loss nonlinear paths (parameterized as Bézier curves) in weight space, even when linear interpolation between them encounters high-loss barriers. This finding established that the loss landscape contains simple, curved corridors connecting distinct modes, motivating subsequent work on identifying conditions under which even simpler linear interpolation suffices. Frankle et al. (2020) later formalized the notion of linear mode connectivity, showing that models sharing a common training trajectory, particularly those fine-tuned from a shared initialization, can be linearly interpolated without traversing high-loss barriers, a finding with important implications for model merging strategies that rely on simple averaging operations.
@@
Loss surface geometry also has direct implications for generalization after merging. Flat minima, characterized by low curvature in the Hessian eigenspectrum, have been associated with improved generalization properties (Izmailov et al., 2018). When merging models that reside in flat regions of the loss landscape, the resulting averaged parameters are more likely to remain within a low-loss region, thereby preserving generalization capabilities. Conversely, sharp minima (Keskar et al., 2017) with high curvature present challenges for merging, as even small perturbations induced by averaging can push parameters into high-loss regions. Lubana et al. (2023) reinforce this from a mechanistic perspective, showing that low-loss linear connectivity between two models implies similarity in their underlying computational mechanisms; models relying on fundamentally different features cannot be connected by simple interpolation paths.
@@
Mathematical analysis of overparameterized networks reveals that the loss surface increasingly resembles a high-dimensional valley structure as model capacity grows. Let $\mathcal{M}$ denote the manifold of global minima; in the overparameterized regime, $\mathcal{M}$ becomes a connected, often convex subspace within which arbitrary linear combinations of solutions remain optimal. This geometric property, combined with rigorous proofs that wide networks trained with SGD are linearly mode connected (Ferbach et al., 2024), provides theoretical justification for why weight averaging succeeds when models share sufficient structural commonality, since their solutions lie within the same connected low-loss region, rendering interpolation a geometrically valid operation.
@@ 2.2 Linear Mode Connectivity Theory and Basin Characteristics
Mode connectivity, the existence of low-loss paths connecting different trained solutions in weight space, provides the central theoretical lens for understanding when and why model merging succeeds. Building on the geometric insights of the preceding section, this framework formalizes the conditions under which interpolation between independently trained solutions remains viable. This concept has become central to explaining the empirical success of weight interpolation methods, as thoroughly surveyed in (Yang et al., 2024b).
@@
Formally, two neural network solutions $\theta_{1}$ and $\theta_{2}$ are said to exhibit linear mode connectivity (Frankle et al., 2020) if the loss along the linear interpolation path $\theta(\alpha)=(1-\alpha)\theta_{1}+\alpha\theta_{2}$ for $\alpha\in[0,1]$ satisfies $\mathcal{L}(\theta(\alpha))\leq\max(\mathcal{L}(\theta_{1}),\mathcal{L}(\theta_{2}))+\epsilon$ for some small tolerance $\epsilon\geq 0$. When $\epsilon=0$, the models reside in the same loss basin with no interpolation barrier. More generally, nonlinear mode connectivity (Garipov et al., 2018; Draxler et al., 2018) relaxes the path constraint to allow curved trajectories, typically parameterized as Bézier curves or piecewise linear paths. This distinction has practical implications, as linear mode connectivity (Frankle et al., 2020) enables direct weight averaging, while nonlinear connectivity may require more involved merging procedures.
@@
Analysis of loss barriers, the maximum loss increase encountered along interpolation paths, provides quantitative measures for characterizing mode connectivity strength. Singh et al. (2024) present a systematization of methods for computing and interpreting these barriers, noting that barrier height correlates inversely with merging success. Specifically, given solutions $\theta_{1}$ and $\theta_{2}$, the loss barrier is defined as follows:
$\Delta\mathcal{L}=\max_{\alpha\in[0,1]}\mathcal{L}(\theta(\alpha))-\max(\mathcal{L}(\theta_{1}),\mathcal{L}(\theta_{2})).$ (1)
@@
When $\Delta\mathcal{L}\leq\epsilon$ for some small threshold $\epsilon$, we say the models exhibit $\epsilon$-linear mode connectivity (Frankle et al., 2020). The relationship between loss barrier magnitude and merging success can be formally characterized through the following theorem.
Proposition 1 (Merging Error Bound).
@@
Let $\theta_{1},\theta_{2}$ be two models with shared pretrained initialization $\theta_{0}$, and let $\theta_{\text{avg}}=\frac{1}{2}(\theta_{1}+\theta_{2})$. Under $L$-smoothness of the loss function $\mathcal{L}$ and the assumption that both models are at approximate stationary points ($\|\nabla\mathcal{L}(\theta_{i})\|\approx 0$, as is typical for trained models), a standard result from convex analysis yields the following bound on the performance degradation of the averaged model (see, e.g., Singh et al. 2024 for a refined, Hessian-dependent generalization):
$\mathcal{L}(\theta_{\text{avg}})-\frac{1}{2}(\mathcal{L}(\theta_{1})+\mathcal{L}(\theta_{2}))\leq\frac{L}{8}\|\theta_{1}-\theta_{2}\|^{2}.$ (2)
@@
This bound reveals why shared initialization is critical, since models fine-tuned from the same pretrained checkpoint exhibit smaller $\|\theta_{1}-\theta_{2}\|$, yielding tighter merging guarantees.
@@
Shared pretrained initialization is the single most important factor enabling strong mode connectivity in the LLM era. When multiple models are fine-tuned from a common pretrained checkpoint, they exhibit considerably stronger linear mode connectivity (Frankle et al., 2020) compared to independently trained models (Garipov et al., 2018). Fine-tuning typically induces relatively small parameter perturbations that remain within the same loss basin as the pretrained model, whereas training from random initialization explores disconnected regions of the loss surface.
@@
Empirical investigations have used various path parameterizations to probe connectivity structure. Quadratic Bézier curves, parameterized as $\theta(\alpha)=(1-\alpha)^{2}\theta_{1}+2\alpha(1-\alpha)\theta_{m}+\alpha^{2}\theta_{2}$ with learnable midpoint $\theta_{m}$, can discover low-loss nonlinear paths even when linear barriers exist. Ren et al. (2025) revisited mode connectivity through the lens of Bézier surfaces, providing richer characterizations of the loss landscape structure between trained solutions beyond simple linear paths.
@@
Prior work on activation renormalization (Jordan et al., 2023) revealed that apparent loss barriers often arise from feature distribution mismatches rather than genuine geometric separation, proposing renormalization procedures to restore connectivity. This finding suggests that many observed failures of linear mode connectivity (Frankle et al., 2020) stem from correctable statistical artifacts rather than core incompatibility between solutions.
@@
Understanding basin characteristics extends beyond barrier analysis to encompass basin width and shape. As shown by Lubana et al. (2023), the success of interpolation between two models depends not only on loss barrier height but also on the degree to which the models share underlying computational mechanisms. These theoretical insights establish that shared pretraining serves as a powerful mechanism for ensuring models occupy connected regions amenable to linear combination. However, even when models theoretically reside in connected basins, the success of direct weight averaging depends critically on how the models are represented, specifically whether their hidden units are aligned, a consideration governed by the symmetries inherent in neural network weight spaces.